Hi,
find all solutions of
cos^2x-3 cosx = 4 on[0, 2pi)
a)pi
b)pi/2, 3pi/2
c)0, pi, 2pi
d)pi/4,3pi/4,5pi/4,7pi/4

Answer by lwsshak3(11628)

find all solutions of
cos^2x-3 cosx = 4 on[0, 2pi)
cos^2x-3cosx-4=0
(cosx-4)(cosx+1)=0
cosx≠4
cosx=-1
x=π (ans. a)